OpenStudy (anonymous):

Solve the equation 2x+1 = 30. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

*2^x+1+30 so the x+1 are exponents

OpenStudy (anonymous):

2^(x+1)=30 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2^(x+1)=30 ln(2^(x+1))=ln(30) (x+1)(ln2)=ln(30) x+1=ln(30)/ln(2) x= 4.9-1 approx. x=3.9 approx